# Addendum to “Power-law decay of the fraction of the mixed eigenstates in kicked top model with mixed-type classical phase space”

### v2

@@ Abstract
By using the Krylov subspace technique to generate the spin coherent states in kicked top model, a prototype model for studying quantum chaos, the accessible system size for studying the Husimi functions of eigenstates can be much larger than that reported in the literature and our previous study Phys. Rev. E 108, 054217 (2023). In the fully chaotic kicked top, we find that the mean Wehrl entropy localization measure approaches the prediction given by the Circular Unitary Ensemble. In the mixed-type case, we identify mixed eigenstates by the overlap of the Husimi function with regular and chaotic regions in classical compact phase space. Numerically, we show that the fraction of mixed eigenstates scales as $j^{-\zeta}$, a power-law decay as the system size $j$ increases, across nearly two orders of magnitude. This provides supporting evidence for the principle of uniform semiclassical condensation of Husimi functions and the Berry-Robnik picture in the semiclassical limit.
@@ I Introduction
In this addendum paper, we use the Krylov subspace techniques [27, 28, 29, 30] to generate the SCS and avoid the factorials. This allows for a significantly larger accessible system size for studying Husimi functions, compared to that reported in literature and our previous study. In the fully chaotic QKT, we find that the mean Wehrl entropy localization measure (ELM) approaches the prediction given by the Circular Unitary Ensemble (CUE). In the mixed-type case, we show a power-law decrease of the fraction of mixed eigenstates as system size increases over nearly two orders of magnitude, corroborating the PUSC and the Berry-Robnik picture in the semiclassical limit.
@@ II Quantum kicked top
where the dynamical variables of the top are three components of the angular momentum operators of the spin-$j$ system, can also be expressed in terms of $2j$ collective spin-$\frac{1}{2}$ Pauli operators, for example $J_{z}=\sum_{k=1}^{2j}\sigma_{z}^{(k)}/2$. The dimension of the Hilbert space is $N=2j+1$, and the squared angular momentum is conserved, $J^{2}=j(j+1)$ with $j$ integer or half-integer. The first term in Eq. (1) describes a precessional rotation about the $x$-axis with angular frequency $\alpha$, the second term denotes a torsional rotation around the $z$-axis with strength $\gamma$ (with the period set to unity).
@@
where $|j,m,\pm\rangle=(|j,m\rangle\pm|j,-m\rangle)/\sqrt{2}$, $1\leq m\leq j$ for the odd parity, and $0\leq m\leq j$ for the even parity where $|j,0,+\rangle$ is not normalized. It should be noted that at $\alpha=\pi/2$, the QKT has a further symmetry [25]. In this addendum paper, we set $\alpha=11\pi/19$, the same as the previous work.
@@ II.1 The transition to Hamiltonian chaos
where $\tilde{x}_{n}=x_{n}$, $\tilde{y}_{n}=y_{n}\cos\alpha-z_{n}\sin\alpha$, and $\tilde{z}_{n}=y_{n}\sin\alpha+z_{n}\cos\alpha$. This classical map $\mathbf{x}_{n+1}=f(\mathbf{x}_{n})$ is restricted on the unit sphere $\mathbb{S}^{2}$ and the Jacobian matrix $\partial f/\partial\mathbf{x}_{n}=\Omega_{n}\cdot R_{\alpha}$, where
@@
The stroboscopic dynamics of the classical kicked top (CKT) is symplectic since $\det(\partial f/\partial\mathbf{x}_{n})=1$, and $\mathbb{S}^{2}$ as the phase space of the classical top can be parametrized as: $x=\sin\theta\cos\phi$, $y=\sin\theta\sin\phi$, $z=\cos\theta$, with $z$ and $\phi$ being the canonical pair. This map undergoes a transition from integrability to chaos with increasing kicking strength $\gamma$, as illustrated in Fig. 1(a), the Poincaré section of the classical map, for different values of $\gamma$. In the classical phase space, regular orbits predominate at small values of $\gamma$. As the kicking strength increases, mixed dynamics emerge, characterized by the coexistence of regular orbits and chaotic motion. Eventually, for larger $\gamma$, the system transitions into full chaos, as illustrated in the rightmost panel of Fig. 1(a).
To qualitatively demonstrate the transition to chaos in the classical map, in this addendum, we use the smaller alignment index (SALI) [31, 32, 33] which rely on the evolution of deviation vectors from a given orbit, to detect regular and chaotic motion of the symplectic map. For chaotic orbits, it was proven that SALI$(t)\propto e^{(-L_{1}-L_{2})t}$ [34], with $L_{1},L_{2}$ being the two largest Lyapunov exponents (LEs). While in three-dimensional symplectic map $L_{2}=0$, therefore for CKT, the SALI is directly related to the largest LE as SALI$(t)\propto e^{-L_{1}t}$. Fig. 1(b) illustrates the logarithmic SALI values after 300 kicks for grid points, at each point plotted using an assigned color according to its value. Upon comparing with the corresponding Poincaré sections in Fig. 1(a), in addition to the resemblance, we also acquire a much more detailed picture of the regions where chaotic or regular motion occurs, and at the borders between these regions we find intermediate colors which correspond to sticky orbits.
@@ III.1 Spin coherent state and Husimi functions
Numerical studies of Husimi functions in the past usually employed Eq. (11b) for the SCS representation of quantum states. However, the presence of large factorials in this equation imposes limitations on the attainable values of $j$, typically only a few hundred. A further examination of Eq. (11a) shows that the SCS can be obtained directly from a unitary transformation of the polar state $|\theta=0,\phi\rangle=|j,j\rangle$. This unitary transformation is an exponential of a sparse matrix, because in the Dicke basis
@@
Exponential of sparse matrix acting on vectors can be effectively computed using Krylov subspace technique (see Ref. [28]), this alternative approach bypasses the problem of large factorials and enables the attainment of significantly larger value of $j$ (differing by several orders of magnitude, depending on the sparsity), compared with the previous studies. One can take an initial look at the Husimi functions of eigenstates provided in Sec. IV.1 and more illustrations in the Appendix, where $j=2^{12}$.
@@ III.2 Localization of eigenstates
Figure 3: Mean Shannon ELM $\langle L\rangle$ from the odd parity in two different basis and mean Wehrl ELM $\langle\mathcal{L}\rangle$, versus $\gamma$ for QKT with $j=2^{12}$ and $\alpha=11\pi/19$. The dashed horizontal lines indicate the mean Shannon ELM of COE and Wehrl/Shannon ELM of CUE.
In Fig. 3, we study the mean Shannon ELM $\langle L\rangle$ in the basis of $J_{x,-}$ and the odd basis $|j,m,-\rangle$, for eigenstates with odd parity (we found for the even parity, it is the same), as well as the mean Wehrl ELM $\langle\mathcal{L}\rangle$. It shows that the Shannon ELM in the basis of $J_{x,-}$ saturates to the COE value $\langle L\rangle_{R}$ for $\gamma\gtrsim 4.4$, and before that, it exhibits two different types of growth: for $\gamma>2$, it grows faster than for $\gamma\leq 2$. While $\langle L\rangle$ in $|j,m,-\rangle$ basis also eventually reaches the COE value $\langle L\rangle_{R}$ for $\gamma\gtrsim 4.4$, but the growth pattern leading up to saturation is markedly different and does not follow the classical transition route to chaos (as shown in Fig. 2(a), and see Appendix A for frurther discussions). This difference is clearly due to the basis-dependent of Shannon entropy. The Wehrl ELM is calculated by using a grid of $N_{p}=200\times 400$ points in the $(\theta,\phi)$ space, with each point holding a coherent state, where all the Husimi functions are normalized according to the discretization of Eq. (14), as
@@
The transition of mean Wehrl ELM versus $\gamma$ follows exactly the classical route. For $\gamma\leq 2$, it maintains a value equal to $\langle\mathcal{L}\rangle_{J_{x}}\sim N^{-1/2}$, representing the mean Wehrl ELM of the operator $J_{x}$ (see numerical result shown in Fig. 11 and proof in Appendix B), which is close to 0 for large $N$. It then increases until it saturates for $\gamma\gtrsim 4.4$, to a value close to $\langle\mathcal{L}\rangle_{R}$ from CUE.
We have examined $\langle\mathcal{L}\rangle$ with increasing number of grid points to $N_{p}=250\times 500$, for $j=2^{12}$. In Fig. 4 it shows that the mean Wehrl ELM is nearly unchanged versus the increasing number of grid points, for three different $\gamma$. Numerically, a larger $j$ requires more grid points, because the number of Planck cells is $N_{c}\sim\hbar_{eff}^{-1}=j$. In this work, for the following calculations, we set $N_{p}=200\times 400\gg j$, for the subsequent calculations to ensure the convergence of all numerical results. Our findings indicate that the mean Wehrl entropy of fully chaotic QKT is slightly smaller than that $\langle S\rangle_{\mathcal{R}}$ from CUE, and it was conjectured in Ref. [50] that the difference between the mean Wehrl entropy of COE and CUE vanishes in the asymptotic limit $N\to\infty$. However, both a numerical verification at larger $N$ and a formal mathematical proof of this conjecture are still lacking. In this work, using the Krylov subspace technique to generate SCSs, we study the mean Shannon entropy $\langle H\rangle$ and Wehrl entropy $\langle S\rangle$, of the QKT at $\gamma=8$ where the system is fully chaotic, as shown in Fig. 5. The eigenvector statistic reveals that $\langle H\rangle\simeq\ln N-0.7322$, is nearly equal to $\langle H\rangle_{\mathcal{R}}$ of COE. For Shannon ELM, $\langle L\rangle\simeq\langle L\rangle_{\mathcal{R}}$, as shown in the inset (bottom right). On the other hand, the mean Wehrl entropy $\langle S\rangle\simeq\ln N-0.4573$, is close to $\langle S\rangle_{\mathcal{R}}$ of CUE, while the inset in the top left illustrates that $\langle\mathcal{L}\rangle$ approaches the CUE value $\langle\mathcal{L}\rangle_{\mathcal{R}}$ with increasing system size.
@@
In fully chaotic QKT exhibiting COE spectral statistics, the distinct ergodic properties of coherent states in the eigenbasis $|v_{n}\rangle$ and the eigenvectors in the angular momentum basis account for the difference between the CUE result for the mean Wehrl entropy and the COE result for the mean Shannon entropy. Consider chaotic systems without any antiunitary symmtries, described by $N\times N$ CUE random matrices, or systems with one antiunitary symmetry, such as time-reversal, described by COE random matrices. As $N\to\infty$, the components of eigenvector $|\psi_{n}\rangle$ of random matrices, denoted by $x_{i}=|\langle i|\psi_{n}\rangle|^{2}$, follow the $\chi_{\nu}^{2}$ distribution [51]
@@
where $\langle x\rangle$ is the mean value of $\{x_{i}\}$, is equal to $1/N$. For the COE, $\nu=1$, $P_{\nu}(x)$ is the well-known Porter-Thomas distribution [52]. For the CUE, $\nu=2$, and for symplectic ensemble, $\nu=4$. This distribution represents the sum of squares $x=\sum_{i=1}^{\nu}y_{i}^{2}$ if each $y_{i}$ is an independent random variable with a Gaussian probability distribution, with the mean equal to zero and the variance equal to $\sqrt{\langle x\rangle}/\nu$. For large $N$, there are a significant number of small components, so it is more effective to use the distribution of $\{\ln x_{i}\}$ instead of $\{x_{i}\}$, given by [53]
$P_{\nu}($ $\ln x)=\frac{P_{\nu}(\ln x)}{d\ln x/dx}$
@@
Figure 7: (a) A joint distribution $(\mathcal{L},M)$ for 8193 eigenstates of QKT Floquet operator, with $\gamma=2.6$, $\alpha=11\pi/19$ and $j=2^{12}$. Husimi functions of eigenstates isolated by boxes at different positions are plotted from panel (a1) to (d3), the darkest blues show the area where $Q_{n}<10^{-16}$. (b) Logarithmic values of SALI for the corresponding CKT, on the parametrized phase space $(\theta,\phi)$ after 300 kicks.
The expansion coefficients of some fixed vector in the eigenbasis of random matrices will have the same statistical properties. Specifically, the components $|\langle\alpha|\psi_{n}\rangle|^{2}$ of an arbitrary coherent state $|\alpha\rangle$ in the eigenbasis $|\psi_{n}\rangle$ follow a $\chi_{\nu}^{2}$ distribution. In a fully chaotic QKT, the components $Q_{n}(\theta,\phi)$ of the SCS $|\theta,\phi\rangle$ in the eigenbasis $|v_{n}\rangle$ of the Floquet operator should conform to this distribution, $\nu=2$. Similarly, the expansion coefficients $|\langle j,m|v_{n}\rangle|^{2}$ of eigenvectors in the $|j,m\rangle$ basis should follow the distribution, $\nu=1$. In Fig. 6 we show $P(\ln x)$ for both types of expansion coefficients in QKT with $\gamma=8$, which is fully chaotic. It demonstrates that for SCS, $P(\ln Q_{n})$ is closely matches $P_{2}(\ln x)$, while $P(\ln|\langle j,m|v_{n}\rangle|^{2})$ is exactly described by $P_{1}(\ln x)$. An intuitive explanation of this difference is that the expansion coefficients of SCS in the eigenbasis are generally complex numbers, where the real part and the imaginary part are independent Gaussian random numbers.
A more physical argument proposed by Zyczkowski [54] is that there are symmetry lines of the antiunitary time reversal operators $T_{1}$ and $T_{2}$ (defined in Eq. (10)) in the phase space. Coherent states located along these two lines exhibit a $\chi_{1}^{2}$ distribution of COE, while those not follow a $\chi_{2}^{2}$ distribution of CUE. The symmetry lines of $T_{1}$ and $T_{2}$ are given by (See Ref. [25] for details)
@@
a parametrization of these lines on $\mathbb{S}^{2}$ is
@@
In Fig. 6 we show that for SCSs $|\theta_{s},\phi_{s}\rangle$, the distribution of the expansion coefficients $P(Q_{n}(\theta_{s},\phi_{s}))$ exactly follows $\chi_{1}^{2}$ distribution. In a fully chaotic QKT, the SCSs along the symmetry lines (great circle on the surface of sphere), have zero measure in the phase space. Assuming the distribution of $Q_{n}(\theta,\phi)$ is independent of $\theta$ and $\phi$, the mean Wehrl entropy
@@
where $N=2j+1$, this is exactly $\langle S\rangle_{\mathcal{R}}$ of CUE in the asymptotic limit.

### v3

# Further results on the power-law decay of the fraction of the mixed eigenstates in kicked top model with mixed-type classical phase space

@@ Abstract
By using the Krylov subspace technique to generate the spin coherent states in kicked top model, a prototype model for studying quantum chaos, the accessible system size for studying the Husimi functions of eigenstates can be much larger than that reported in the literature and our previous study Phys. Rev. E 108, 054217 (2023). In the fully chaotic kicked top, we show that the mean Wehrl entropy localization measure approaches the prediction given by the Circular Unitary Ensemble, while the spectral statistics follows the Circular Orthogonal Ensemble. For the mixed-type cases, we identify mixed eigenstates by the overlap of the Husimi function with regular and chaotic regions in classical compact phase space. Numerically, we show that the fraction of mixed eigenstates scales as $j^{-\zeta}$, a power-law decay as the system size $j$ increases, across nearly two orders of magnitude, expanding significantly beyond the previously limited range of system sizes. This provides supporting evidence for the principle of uniform semiclassical condensation of Husimi functions and the Berry-Robnik picture in the semiclassical limit.
@@ I Introduction
In this paper, we use the Krylov subspace techniques [27, 28, 29, 30] to generate the SCS and avoid the large factorials. This allows for a significantly larger accessible system size for studying Husimi functions, compared to that reported in literature and our previous study. It is important to note that the issue of large factorials in SCS is similar to the longstanding problem encountered in the numerical evaluation of the Wigner rotation matrix, which is crucial in modern nuclear-structure models. In both cases, directly calculating large factorials at high spins results in significant precision loss, instability, and increased memory consumption [31, 32], as SCS can be derived from a SU(2) rotation. In the fully chaotic QKT, by using SCSs generated through the Krylov subspace method, we find that the mean Wehrl entropy localization measure (ELM) approaches the prediction given by the Circular Unitary Ensemble (CUE), provided that coherent states situated on symmetrical lines exhibit COE eigenvector statistics but have zero measure in phase space, whereas those outside follow CUE eigenvector statistics. For the mixed-type cases, we show a power-law decrease of the fraction of mixed eigenstates as system size increases over nearly two orders of magnitude, corroborating the PUSC and the Berry-Robnik picture in the semiclassical limit.
@@ II Quantum kicked top
where the dynamical variables of the top are three components of the angular momentum operators of the spin-$j$ system, can also be expressed in terms of $2j$ collective spin-$\frac{1}{2}$ Pauli operators, for example $J_{z}=\sum_{k=1}^{2j}\sigma_{z}^{(k)}/2$. The dimension of the Hilbert space is $N=2j+1$, and the squared angular momentum is conserved, $J^{2}=j(j+1)$ with $j$ integer or half-integer (in this work, we consider integer $j$). The first term in Eq. (1) describes a precessional rotation about the $x$-axis with angular frequency $\alpha$, the second term denotes a torsional rotation around the $z$-axis with strength $\gamma$ (with the period set to unity).
@@
where
$|j,m,\pm\rangle=(|j,m\rangle\pm|j,-m\rangle)/\sqrt{2},$ (4)
with $1\leq m\leq j$ for the odd parity, and $0\leq m\leq j$ for the even parity where $|j,0,+\rangle$ is not normalized. It should be noted that at $\alpha=\pi/2$, the QKT has a further symmetry [25]. In this paper, we set $\alpha=11\pi/19$, the same as the previous work.
@@ II.1 The transition to Hamiltonian chaos
where $\tilde{x}_{n}=x_{n}$, $\tilde{y}_{n}=y_{n}\cos\alpha-z_{n}\sin\alpha$, and $\tilde{z}_{n}=y_{n}\sin\alpha+z_{n}\cos\alpha$. This classical map $\mathbf{x}_{n+1}=F_{c}(\mathbf{x}_{n})$ is restricted on the unit sphere $\mathbb{S}^{2}$ and the Jacobian matrix $\partial F_{c}/\partial\mathbf{x}_{n}=\Omega_{n}\cdot R_{\alpha}$, where
@@
The stroboscopic dynamics of the classical kicked top (CKT) is symplectic since $\det(\partial F_{c}/\partial\mathbf{x}_{n})=1$, and $\mathbb{S}^{2}$ as the phase space of the classical top can be parametrized as: $x=\sin\theta\cos\phi$, $y=\sin\theta\sin\phi$, $z=\cos\theta$, with $z$ and $\phi$ being the canonical pair. Two symmetries of CKT can be found by two involutions [25], defined by
$T_{1}^{c}(\mathbf{x})$ $=(x,-y\cos\alpha+z\sin\alpha,y\sin\alpha+z\cos\alpha),$ (8a)
$T_{2}^{c}(\mathbf{x})$ $=(x,y\cos\alpha-z\sin\alpha,-y\sin\alpha-z\cos\alpha),$ (8b)
where $(T_{1}^{c})^{2}=(T_{2}^{c})^{2}=1$ with the determinants $\det(\partial T_{1}^{c}/\partial\mathbf{x})=\det(\partial T_{2}^{c}/\partial\mathbf{x})=-1$. These two involutions yield time reversal operations for the classical map $F_{c}$ in the sense that $T_{1}^{c}F_{c}T_{1}^{c}=T_{2}^{c}F_{c}T_{2}^{c}=F_{c}^{-1}$, indicating that the map has double reversible dynamics.
CKT undergoes a transition from integrability to chaos with increasing kicking strength $\gamma$, as illustrated in Fig. 1(a), the Poincaré section of the classical map, for different values of $\gamma$. In the classical phase space, regular orbits predominate at small values of $\gamma$. As the kicking strength increases, mixed dynamics emerge, characterized by the coexistence of regular orbits and chaotic motion. Eventually, for larger $\gamma$, the system transitions into full chaos, as illustrated in the rightmost panel of Fig. 1(a).
To qualitatively demonstrate the transition to chaos in the classical map, in this work, we use the smaller alignment index (SALI) [33, 34, 35] which rely on the evolution of deviation vectors from a given orbit, to detect regular and chaotic motion of the symplectic map. For chaotic orbits, it was proven that SALI$(t)\propto e^{(-L_{1}-L_{2})t}$ [36], with $L_{1},L_{2}$ being the two largest Lyapunov exponents (LEs). While in three-dimensional symplectic map $L_{2}=0$, therefore for CKT, the SALI is directly related to the largest LE as SALI$(t)\propto e^{-L_{1}t}$. Fig. 1(b) illustrates the logarithmic SALI values after 300 kicks for grid points, at each point plotted using an assigned color according to its value. Upon comparing with the corresponding Poincaré sections in Fig. 1(a), in addition to the resemblance, we also acquire a much more detailed picture of the regions where chaotic or regular motion occurs, and at the borders between these regions we find intermediate colors which correspond to sticky orbits.
Fig. 2(a) shows a magnified SALI plot of the phase space, for the CKT at $\gamma=4$, focusing on the region around the regular island depicted in the middle panel of Fig. 1 (b). Clearly, smaller islands of regular motion appear around the main island, organized in a hierarchical way [3]. In addition, along the border between the regular islands and the chaotic sea, the existence of stickiness is evident, indicated by intermediate SALI values. In deterministic dynamical systems, stickiness is associated with the partial barriers, such as remnants of destroyed invariant tori [37], which slow down the transport between different regions of phase space. Fig. 2(b) illustrates the typical time evolution of the SALI for three types of orbits: regular, sticky, and chaotic. It shows that the sticky orbit remains near the border between the chaotic and regular region for a very long time, requiring significantly more iterations than typical chaotic orbits to reveal its true chaotic behavior.
Figure 2: (a) Magnified view of the SALI (after 300 kicks) plot for the CKT with $\gamma=4$, highlighting the left-bottom region around the regular island in the middle panel of Fig. 1(b). (b) Evolution of the SALI with respect to time t, for three initial conditions chosen from panel (a), denoted by R (regular), S (sticky) and C (chaotic). The SALI of sticky orbit takes a longer time to approach the threshold value $10^{-8}$, as it remains near the border between the chaotic and regular regions up to approximately $t=200$.
@@ III.1 Spin coherent state and Husimi functions
Numerical studies of Husimi functions in the past usually employed Eq. (13b) for the SCS representation of quantum states. However, the presence of large factorials in this equation imposes limitations on the attainable values of $j$, typically only a few hundred. This limitation is not only due to the memory consumption and computational complexity associated with the direct calculation of large factorials (such as $n!$ for $n>1000$), as it also stems from the numerical instability and significant precision loss. This issue is reminiscent of the longstanding problem in the numerical evaluation of the Wigner rotation matrix [31], as for SCSs $|\theta,0\rangle=\exp(-i\theta J_{y})|j,j\rangle=\sum_{m}d_{mj}^{j}|j,m\rangle$, where the expansion coefficients are given by the Wigner $d$-matrix.
A further examination of Eq. (13a) shows that the SCS can be obtained directly from a unitary transformation of the polar state $|\theta=0,\phi\rangle=|j,j\rangle$. This unitary transformation is an exponential of a sparse matrix, because in the Dicke basis
@@
Exponential of sparse matrix acting on vectors can be effectively computed using Krylov subspace technique (see Ref. [28]). This alternative approach bypasses the problem of large factorials and enables the attainment of significantly larger value of $j$ (differing by several orders of magnitude, depending on the sparsity), compared with the previous studies. One can take an initial look at the Husimi functions of eigenstates provided in Sec. IV.1 and more illustrations in the Appendix, where $j=2^{12}$. The Krylov subspace technique is a well-known method in numerical linear algebra and can also be applied to the numerical calculation of Wigner rotation matrices for higher spins, for comparison with other effective algorithm [32].
@@ III.2 Localization of eigenstates
In Fig. 4, we study the mean Shannon ELM $\langle L\rangle$ in the eigenbasis of $J_{x,-}$ and the basis $|j,m,-\rangle$ for odd parity defined in Eq. (4), for eigenstates with odd parity (we found for the even parity, it is the same), as well as the mean Wehrl ELM $\langle\mathcal{L}\rangle$. It shows that the Shannon ELM in the basis of $J_{x,-}$ saturates to the COE value $\langle L\rangle_{R}$ for $\gamma\gtrsim 4.4$, and before that, it exhibits two different types of growth: for $\gamma>2$, it grows faster than for $\gamma\leq 2$. While $\langle L\rangle$ in $|j,m,-\rangle$ basis also eventually reaches the COE value $\langle L\rangle_{R}$ for $\gamma\gtrsim 4.4$, but the growth pattern leading up to saturation is markedly different and does not follow the classical transition route to chaos (as shown in Fig. 3(a), and see Appendix A for further discussions). This difference is clearly due to the basis-dependence of Shannon entropy. The Wehrl ELM is calculated by using a grid of $N_{p}=200\times 400$ points in the $(\theta,\phi)$ space, with each point holding a coherent state, where all the Husimi functions are normalized according to the discretization of Eq. (16), as
@@
The transition of mean Wehrl ELM versus $\gamma$ follows exactly the classical route. For $\gamma\leq 2$, it maintains a value equal to $\langle\mathcal{L}\rangle_{J_{x}}\sim N^{-1/2}$, representing the mean Wehrl ELM of the operator $J_{x}$ in the basis of unperturbed Hamiltonian (see proof in Appendix B), which is close to 0 for large $N$. It then increases until it saturates for $\gamma\gtrsim 4.4$, to a value close to $\langle\mathcal{L}\rangle_{R}$ from CUE.
Figure 4: Mean Shannon ELM $\langle L\rangle$ calculated for the odd parity, in the eigenbasis of $J_{x,-}$ and the basis $|j,m,-\rangle$ for odd parity defined in Eq. (4) and mean Wehrl ELM $\langle\mathcal{L}\rangle$, versus $\gamma$ for QKT with $j=2^{12}$ and $\alpha=11\pi/19$. The dashed horizontal lines indicate the mean Shannon ELM from COE statistics and Wehrl/Shannon ELM of CUE.
Numerically, a larger $j$ requires more grid points, because the number of Planck cells is $N_{c}\sim\hbar_{eff}^{-1}=j$. In this work, for the following calculations, we set $N_{p}=200\times 400\gg j$, for the subsequent calculations to ensure the convergence of all numerical results (see details about the numerical convergence in Appendix B). Our findings indicate that the mean Wehrl entropy of fully chaotic QKT is slightly smaller than that $\langle S\rangle_{\mathcal{R}}$ from CUE, and it was conjectured in Ref. [53] that the difference between the mean Wehrl entropy of COE and CUE vanishes in the asymptotic limit $N\to\infty$. However, both a numerical verification at larger $N$ and a formal mathematical proof of this conjecture are still lacking. In this work, using the Krylov subspace technique to generate SCSs, we study the mean Shannon entropy $\langle H\rangle$ and Wehrl entropy $\langle S\rangle$, of the QKT at $\gamma=8$ where the system is fully chaotic, as shown in Fig. 5. The eigenvector statistic reveals that $\langle H\rangle\simeq\ln N-0.7322$, is nearly equal to $\langle H\rangle_{\mathcal{R}}$ of COE. For Shannon ELM, $\langle L\rangle\simeq\langle L\rangle_{\mathcal{R}}$, as shown in the inset (bottom right). On the other hand, the mean Wehrl entropy $\langle S\rangle\simeq\ln N-0.4573$, is close to $\langle S\rangle_{\mathcal{R}}$ of CUE, while the inset in the top left shows that $\langle\mathcal{L}\rangle$ approaches the CUE value $\langle\mathcal{L}\rangle_{\mathcal{R}}$ with increasing system size.
@@
In fully chaotic QKT exhibiting COE spectral statistics, the distinct ergodic properties of coherent states in the eigenbasis $|v_{n}\rangle$ and the eigenvectors in the angular momentum basis account for the difference between the CUE result for the mean Wehrl entropy and the COE result for the mean Shannon entropy. Consider chaotic systems without any antiunitary symmetries, described by $N\times N$ CUE random matrices, or systems with one antiunitary symmetry, such as time-reversal, described by COE random matrices. As $N\to\infty$, the components of eigenvector $|\psi_{n}\rangle$ of random matrices, denoted by $x_{i}=|\langle i|\psi_{n}\rangle|^{2}$, follow the $\chi_{\nu}^{2}$ distribution [54]
@@
where $\langle x\rangle$ is the mean value of $\{x_{i}\}$, equal to $1/N$. For the COE, $\nu=1$, $P_{\nu}(x)$ is the well-known Porter-Thomas distribution [55]. For the CUE, $\nu=2$, and for symplectic ensemble, $\nu=4$. This distribution represents the sum of squares $x=\sum_{i=1}^{\nu}y_{i}^{2}$ if each $y_{i}$ is an independent random variable with a Gaussian probability distribution, with the mean equal to zero and the variance equal to $\sqrt{\langle x\rangle}/\nu$. For large $N$, there are a significant number of small components, so it is more effective to use the distribution of $\{\ln x_{i}\}$ instead of $\{x_{i}\}$, given by [56]
$P_{\nu}($ $\ln x)=\frac{P_{\nu}(x)}{d\ln x/dx}$
@@
The expansion coefficients of some fixed vector in the eigenbasis of random matrices will have the same statistical properties. Specifically, the components $|\langle\alpha|\psi_{n}\rangle|^{2}$ of an arbitrary coherent state $|\alpha\rangle$ in the eigenbasis $|\psi_{n}\rangle$ follow a $\chi_{\nu}^{2}$ distribution. In a fully chaotic QKT, the components $Q_{n}(\theta,\phi)$ of the SCS $|\theta,\phi\rangle$ in the eigenbasis $|v_{n}\rangle$ of the Floquet operator should conform to this distribution, $\nu=2$. Similarly, the expansion coefficients $|\langle j,m|v_{n}\rangle|^{2}$ of eigenvectors in the $|j,m\rangle$ basis should follow the distribution, $\nu=1$. In Fig. 6 we show $P(\ln x)$ for both types of expansion coefficients in QKT with $\gamma=8$, which is fully chaotic. It demonstrates that for all SCSs located on the unit sphere, $P(\ln Q_{n})$ is closely matches $P_{2}(\ln x)$, while $P(\ln|\langle j,m|v_{n}\rangle|^{2})$ is exactly described by $P_{1}(\ln x)$. An intuitive explanation of this difference is that the expansion coefficients of SCS in the eigenbasis are generally complex numbers, where the real part and the imaginary part are independent Gaussian random numbers.
@@
A more physical argument proposed by Zyczkowski [57] suggests that there exists a symmetry line in the phase space and the coherent state located along this line exhibit a $\chi_{1}^{2}$ distribution of COE, while those off the line follow a $\chi_{2}^{2}$ distribution of CUE. Importantly, this single symmetry line has measure zero within the compact phase space. However, instead of a single symmetry line in the phase space, there exists a family of symmetry curves in the classical phase space. These curves are defined as the set of points $\mathbf{x}\in\mathbb{S}^{2}$ that satisfy the condition $\mathcal{I}(\mathbf{x})=\mathbf{x}$, where $\mathcal{I}=T_{1}^{c},T_{2}^{c},T_{1}^{c}F_{c},F_{c}T_{1}^{c},\cdots$ [25, 58]. For instance, the symmetry lines of $T_{1}^{c}$ and $T_{2}^{c}$ are given by
@@
are great circles on the unit sphere, a parametrization of these two lines on $\mathbb{S}^{2}$ is
@@
The symmetry lines for higher-order involutions tend to have more complicated shapes.
In Fig. 6 we show that for SCSs $|\theta_{s},\phi_{s}\rangle$ located on two great circles, the distribution of the expansion coefficients $P(Q_{n}(\theta_{s},\phi_{s}))$ exactly follows $\chi_{1}^{2}$ distribution. We can formally prove that the family of symmetry lines $\{l_{n}\}$, has zero measure in the phase space. In the compact space $\mathbb{S}^{2}$, by defining the Lebesgue outer measure $m^{*}$ for each $l_{n}$, it is evident each line on $\mathbb{S}^{2}$ has $m^{*}(l_{n})=0$. Considering all the involutions in $\mathcal{I}$, one can establish a bijection of each involution and the natural number, proving that $\{l_{n}\}$ is countable with $n\in\mathbb{N}$. From the weaker property of countable subadditivity [59, 60], we see
$m^{*}(\bigcup_{n\in\mathbb{N}}l_{n})\leq\sum_{n\in\mathbb{N}}m^{*}(l_{n})=0,$ (26)
meaning that the symmetry lines are of zero measure. Assuming the distribution of $Q_{n}(\theta,\phi)$ is independent of $\theta$ and $\phi$, for the mean Wehrl entropy, we have
@@
where $N=2j+1$, which is exactly $\langle S\rangle_{\mathcal{R}}$ of CUE in the asymptotic limit.
Figure 7: (a) A joint distribution $(\mathcal{L},M)$ for 8193 eigenstates of QKT Floquet operator, with $\gamma=2.6$, $\alpha=11\pi/19$ and $j=2^{12}$. States cluster at the overlap index extremes—regular at one end, chaotic at the other—with mixed states sparsely in between. Husimi functions of eigenstates isolated by boxes at different positions are plotted from panel (a1) to (d3), the darkest blues show the area where $Q_{n}<10^{-16}$. (b) Logarithmic values of SALI for the corresponding CKT, on the parametrized phase space $(\theta,\phi)$ after 300 kicks.